OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10 Answers x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (anonymous):

do u know what || means?

OpenStudy (anonymous):

im sry if i was rude :/

OpenStudy (anonymous):

Do you want help figuring out how to do it, or just the answer?

OpenStudy (anonymous):

@noneedtoknow either way u need to help figure out how to do it :) :P , answer eventually comes :)

OpenStudy (anonymous):

Theres no solution

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

the actually is a solution...

OpenStudy (anonymous):

I was not sure if I was right so I post on here to see what others got

OpenStudy (anonymous):

@GotToLoveJenny there's a solution!

OpenStudy (anonymous):

actually solutions :)

OpenStudy (anonymous):

@Espree14 , give me ur reasoning on this question, what u got till now?

OpenStudy (anonymous):

x = −3 and x = 1

OpenStudy (anonymous):

Oh if it the one above than yes their is can somebody show me how to get it

OpenStudy (anonymous):

But there isnt if there is a solution prove it please

OpenStudy (anonymous):

|| is called modulus anything in between those two Pipes( | ) will become positive for example | -2 | = 2 | 2 | = 2 so by this if we have an equation |x| = 2 then x maybe 2 or -2

OpenStudy (anonymous):

yes okay....keep going

OpenStudy (anonymous):

so for modulus problems we always have to take two cases * one positive * one negative so here −5|x + 1| = 10 lets just simplify before we split our modulus, ie −|x + 1| = 2 => |x + 1| = -2 right?

OpenStudy (anonymous):

Yes I think so

OpenStudy (anonymous):

im going to check my work

OpenStudy (anonymous):

now here we need to take two cases (x+1) which is positive and -(x+1) which is negative as |x+1| = | -(x+1) | so now x+1 = -2 => x = -3 -(x+1) = -2 => x+1 = 2 => x = 1

OpenStudy (anonymous):

i still got no solution

OpenStudy (anonymous):

Nevermind, none of these options could leave you with a positive 10.. my bad, you where right @GotToLoveJenny!

OpenStudy (anonymous):

yayyyyyyyyyyyyyyyyyyyyyyyyy

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

OH MY BAD!, @GotToLoveJenny you are right! sorry im extremely sorry

OpenStudy (anonymous):

lol no its okay your perfectly fine

OpenStudy (anonymous):

@SandeepReddy